Science and medicine. Celsius temperature readings can be converted to
Fahrenheit readings using the formula E=9/5C+32. What is the Fahrenheit
temperature that corresponds to each of the following Celsius temperatures: -10,
0, 15, 100?

Found 2 solutions by checkley71, sarahschober:
Answer by checkley71(8403)

E=9/5C+32
E=9/5*10+32
E=18+32=50
--------------------
E=9/5*0+32=32
----------------------
E=9/5*15+32
P=27+32=59
---------------------
P=9/5*100+32
P=180+32=212

Answer by sarahschober(35)

-10=14
0=32
15=59
100=212